OpenStudy (askswatw260):

The height above the ground of a ball thrown up with velocity of 96 feet per second from a height of 6 feet is 6+96t-16t^2 feet, where t is the time in seconds. According to this model, how high is the ball after 7 seconds? Explain. I substituted in for t, but I got a negative answer.

OpenStudy (anonymous):

heightt = 6+ 96t - 16t^2 so if t = 7 heightt = 6+ 96t - 16t^2 heightt = 6+ 96(x7) - 16(x7^2) heightt = 6+ 96(x7) - 16(x49) heightt = ??? @Askswatw260 ?

OpenStudy (askswatw260):

what do you mean???? @Jack4

OpenStudy (whpalmer4):

No, you plug in t = 7, not t = x7 \[h(t) = 6+96t - 16t^2\]\[h(7) = 6+96(7)-16(7)^2 = \]

OpenStudy (anonymous):

-106, right?

OpenStudy (askswatw260):

yeah I got that but how is that so? @Jack4

OpenStudy (whpalmer4):

yes, -106 is the number returned. A plot of the function from t = 0 to t = 7 may help understand what is going on here...

OpenStudy (askswatw260):

for the worksheet how would I explain -106?

OpenStudy (whpalmer4):

Well, first, do you understand what the 3 terms in that equation represent?

OpenStudy (askswatw260):

I pretty sure I do

OpenStudy (anonymous):

it's because you consider the speed it's thrown at to be positive... have a look at drawing to follow

OpenStudy (askswatw260):

thank you @Jack4 and @whpalmer4

OpenStudy (anonymous):



OpenStudy (anonymous):

all good dude

OpenStudy (whpalmer4):

The reason you get a negative number is that the equation doesn't have any knowledge of where the ground actually is, and that the ball will stop there. t=7 seconds is a time beyond the time of impact. The equation is a model which only applies to the domain \(t = 0 \text{ to }t\approx 6.062\)